OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

Okay, 2x + 3y = 1200 2x + 3y = 1200 Slope intercept form is y = mx + b, where 'm' is the slope, and 'b' is the y-intercept. So first we subtract '2x' to both sides: 3y = -2x + 1200 Now we divide 3 to both sides: y = -2/3x + 400 So this is slope-intercept form, what are the slope and y-intercept here?

OpenStudy (anonymous):

The y = -2/3+400 & Its A 6 part question :(

OpenStudy (anonymous):

I know, I've answered this question like 2 times in the past.

OpenStudy (anonymous):

Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area.

OpenStudy (anonymous):

y = -2/3x + 400 This is the equation in slope-intercept form, can you tell me what the slope and y-intercept are? It's in the form of: y = mx + b Where 'm' is the slope, and 'b' is the y-intercept.

OpenStudy (anonymous):

Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area.

OpenStudy (anonymous):

Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (anonymous):

y = -2/3x + 400 This is the equation in slope-intercept form, can you tell me what the slope and y-intercept are? It's in the form of: y = mx + b Where 'm' is the slope, and 'b' is the y-intercept.

OpenStudy (anonymous):

That's It @iGreen.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Tell me the slope and y-intercept in y = -2/3x + 400

OpenStudy (anonymous):

the -2/3x +400

OpenStudy (anonymous):

Yes, the slope will be -2/3..and the y-intercept will be 400.

OpenStudy (anonymous):

So if we are to graph this, we can first plot the y-intercept, then draw a line that has -2/3 slope through it.

OpenStudy (anonymous):

Okay .

OpenStudy (anonymous):

That takes care of this: "Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences."

OpenStudy (anonymous):

"Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences." Function notation is just replacing 'y' with 'f(x)' y = -2/3x + 400 So replace 'y' with 'f(x)'

OpenStudy (anonymous):

so it would be f(x) - 2/3 +400?

OpenStudy (anonymous):

f(x) = -2/3x + 400 Yes, but we keep the equal sign.

OpenStudy (anonymous):

Ohh I Forgot Sorry

OpenStudy (anonymous):

"Graph the function using one of the following two options below. On the graph, make sure to label the intercepts." Here's the graph:

OpenStudy (anonymous):

Omg Thank You :)

OpenStudy (anonymous):

what part do we have left ?

OpenStudy (anonymous):

"Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences." We can graph them together, the equation for this one will be: 2x + 3y = 1500. (Note: the red line is the original, the blue line is the new one)

OpenStudy (anonymous):

So they both have the SAME slope, but they have different y-intercepts.

OpenStudy (anonymous):

Okay Ima Make Sure I Put That IN So My Teacher Don't Get Confused.

OpenStudy (anonymous):

"Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations." Okay, first let's find the slope: We can use the 'y' and 'x' intercepts. (0, 300) (450, 0) So plug them into the slope formula: \(m = \dfrac{0-300}{450-0}\) Can you subtract those?

OpenStudy (anonymous):

I Got 50

OpenStudy (anonymous):

I got something different..what's 0 - 300?

OpenStudy (anonymous):

-300

OpenStudy (anonymous):

Yes, and what's 450 - 0?

OpenStudy (anonymous):

450

OpenStudy (anonymous):

Yes, so we have: -300/450 which is the same as -2/3. So our slope is -2/3 and our y-intercept is 300. y = mx + b Where 'm' is the slope, and 'b' is the y-intercept. So we can plug in the slope(-2/3) and the y-intercept(300): y = -2/3x + 300

OpenStudy (anonymous):

So that's our equation..that's all the steps! :D

OpenStudy (anonymous):

1.How can we solve linear equations in one variable? 2. What are the similarities and differences in solving equations and inequalities? 3. What are the steps in solving an absolute value equation? 4. How is solving a literal equation different from solving a one- or two-step equation? This Is It For Now

OpenStudy (anonymous):

Post a new question..we can only have 1 question per post.

OpenStudy (anonymous):

Okay